# On palindromic numerators of bigraded symmetric orbifold Hilbert series and Kostka-Foulkes polynomials

### v2

@@ 1 Introduction
In [1], we introduced and exploited properties of counting formulae for the 1-loop graviton partition function of log gravity [2, 3] on the thermal $\text{AdS}_{3}$ background originally calculated in [4], with the aim of providing a unifying interpretation for both single and multiparticle log sectors. It was shown that the multi-particle states are simply built as a Fock space from products of the single particle states, by reformulating the log partition function in terms of the plethystic exponential (PE), which defines the $n^{\text{th}}$ symmetric product of functions.
Subsequently, we showed in [5] that the moduli space of the log sector is the $S^{n}\left(\mathbb{C}^{2}\right)$ Calabi-Yau orbifold. The modus operandi was to show that the log partition function is a generating function of Hilbert series of the polynomial ring $\mathbb{C}\left[x_{1},\ldots,x_{n},y_{1},\ldots,y_{n}\right]^{S_{n}}$ which is Cohen-Macaulay, and that the numerators of the Hilbert series are palindromic, therefore indicating that the ring being enumerated is Gorenstein, and thus establishing the Calabi-Yau nature of the log sector’s moduli space.
As an expansion in terms of Schur polynomials, i.e Schur symmetric functions expressed as polynomials of power sums, in [6] we showed that the log partition function is a $\tau$-function of the KP hierarchy of soliton evolution equations, using the relationship between Schur functions and infinite-dimensional Grassmann manifold studied in [7]. As such, the free energy in (i.e the logarithm of) the log partition function is a solution of the Kadomtsev-Petviashvili (KP) hierarchy that is isomorphic to an infinite-dimensional Grassmann manifold called the Universal (or Sato’s) Grassmann manifold, and the evolution of the log partition function is parametrized by the motion of a point on that Grassmannian.
@@ 2 Background
The aim of this section is to give an succinct expository account of topics that play a central role in our work. Starting with some definitions of symmetric functions, we provide a brief description of the plethystic substitution notation for symmetric functions, which is essential in subsequently introducing Macdonald polynomials, their associated $q,t$-Kostka polynomials, as well as their geometrical interpretation in terms of bigraded $S_{n}$-representations related to the geometry of the Hilbert scheme of $n$ points in the plane.
@@ 2.1 Symmetric functions
A few well-known natural bases for the ring of symmetric functions [8, 9], all indexed by partitions $\lambda=\left(\lambda_{1},\ldots,\lambda_{k}\right)$ are listed below.
@@
$m_{\lambda}=\sum_{\begin{subarray}{c}i_{a}\neq i_{b}\leavevmode\nobreak\ \forall a,b,\\ i_{a}<i_{b}\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \lambda_{a}=\lambda_{b}\end{subarray}}x_{i_{1}}^{\lambda_{1}}\cdots x_{i_{k}}^{\lambda_{k}}.$ (3)
@@
The irreducible representations of the symmetric group $S_{n}$ are also indexed by partitions $\lambda\vdash n$ [8, 11]. Given an $S_{n}$-module $V$, denote the irreducible $S_{n}$-module by $V_{\lambda}$ such that
@@ 2.3.1 Macdonald polynomials and (q,t)-Kostka Macdonald coefficients
A remarkable new basis for the space of symmetric functions known as the Macdonald polynomials was introduced in [12]. The elements of this basis are denoted $P_{\lambda}\left(X;q,t\right)$ where $\lambda$ is a partition, $X$ is the infinite set of variables $x_{1},x_{2},\ldots$, and $q,t$ are parameters. Their importance stems in part from the fact that by suitable substitutions for the parameters $q$ and $t$, they specialize to many of the well-known bases for the symmetric functions, and one can recover in this manner the Schur functions, the Hall-Littlewood symmetric functions, the Jack symmetric functions, the zonal symmetric functions, the zonal spherical functions, and the elementary and monomial symmetric functions [8].
Given a partition $\lambda$, we define the Young diagram $\mathcal{Y}(\lambda)$ of the partition as the array of lattice squares $\mathcal{Y}(\lambda)=\left\{(i,j)\in\mathbb{N}\times\mathbb{N}:j<\lambda_{i+1}\right\}$, and represent it using the French convention as follows. Given that the parts of $\lambda$ are $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{k}>0$, we let the corresponding Young diagram have $\lambda_{i}$ lattice squares in the $i^{th}$ row (counting from the bottom up), with $n_{\lambda}=\sum_{i=1}^{k}(i-1)\lambda_{i}$, and adopt the Macdonald convention [8] of calling the armarm, legleg, coarmcoarm and colegcoleg of a lattice square $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s$} the parameters $a_{\lambda}$, $l_{\lambda}$, $a^{\prime}_{\lambda}$ and $l^{\prime}_{\lambda}$ giving the number of cells of $\lambda$ that are East, North, West and South of $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s$}, respectively. In this setting, $n_{\lambda}=\sum_{s\in\mathcal{Y}(\lambda)}l(s)$.
[…]
a’ s a a
l’
@@ 2.3.2 Geometry of configuration of $n$ points in the plane
Macdonald conjectured the $K_{\lambda\mu}\left(q,t\right)$ to be polynomials in $q,t$ with positive integer coefficients. In an attempt to prove Macdonald’s conjecture, the so-called modified Macdonald polynomials were introduced in [13] as follows. Let
$H_{\mu}\left(q,t\right)=J_{\mu}\left[\frac{X}{1-t};q,t\right]$ (19)
@@
The remarkable connection between the two-parameter symmetric polynomials discovered in [12], and the geometry of configurations of $n$ points in the affine plane $\mathbb{C}^{2}$ can be summarized as follows. In [13], the authors proposed what became known as the $n!$ conjecture, which states that the transformed Macdonald polynomials $\tilde{H}_{\mu}\left(X;q,t\right)$ could be realized as the bigraded characters of certain modules for the diagonal action of $S_{n}$ on two sets of variables. The $n!$ conjecture was proved in [14], consequently establishing Macdonald positivity, i.e $K_{\lambda\mu}\left(q,t\right)\in\mathbb{N}[q,t]$, by connecting a conjectural representation-theoretic interpretation of the transformed $q,t$-Kostka coefficients to the algebraic geometry of the Hilbert scheme of $n$ points in the plane, $\text{Hilb}^{n}\left(\mathbb{C}^{2}\right)$, which can be described as the algebraic variety parametrizing ideals $I\subseteq\mathbb{C}\left[x,y\right]$, where $\mathbb{C}\left[x,y\right]$ denotes the coordinate ring of the affine plane, such that dim[$_{\mathbb{C}}\mathbb{C}\left[x,y\right]/I=n$ (i.e the quotient $\mathbb{C}\left[x,y\right]/I$ has dimension $n$ as a vector space over $\mathbb{C}$), or equivalently 0-dimensional subschemes of length $n$ in $\mathbb{C}^{2}$ [15, 16].
@@ 2.3.3 The log partition
In [5], we showed that the moduli space of the log sector of log gravity is the orbit variety $\left(\mathbb{C}^{2}\right)^{n}/S_{n}$, where $S_{n}$ acts on $\left(\mathbb{C}^{2}\right)^{n}$ by permuting coordinates. A standard procedure when discussing $S^{n}\left(\mathbb{C}^{2}\right)$ is to start by considering ordered $n$-tuples of points in the plane, which we can denote by $P_{1},\ldots,P_{n}\in\mathbb{C}^{2}$. Assigning the points coordinates $x_{1},y_{1},\ldots x_{n},y_{n}$, the space of all $n$-tuples $\left(P_{1},\ldots,P_{n}\right)$ is identified with $\mathbb{C}^{2n}$. The coordinate ring of $\mathbb{C}^{2n}$ is then the polynomial ring $\mathbb{C}\left[\bf{x,y}\right]=\mathbb{C}\left[x_{1},y_{1},\ldots,x_{n},y_{n}\right]$ in $2n$ variables. It follows from the action of the symmetric group $S_{n}$ on $\mathbb{C}^{2n}$ that the subring of invariants $\mathbb{C}\left[\bf{x,y}\right]^{S_{n}}=\mathbb{C}\left[x_{1},y_{1},\ldots,x_{n},y_{n}\right]^{S_{n}}$ is a direct summand of $\mathbb{C}\left[\bf{x,y}\right]$ as an $S_{n}$-module, and also a $\mathbb{C}\left[\bf{x,y}\right]^{S_{n}}$-module. Furthermore, the Hilbert series of $\mathbb{C}\left[\bf{x,y}\right]^{S_{n}}$ can be computed as a doubly graded ring, by degree in the $\bf{x}$ and $\bf{y}$ variables separately.
@@
Finally, following Macdonald and considering Young diagrams $\mathcal{Y}$ of size $n$, the partition function of the log sector can be expressed as
$Z_{log}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}q^{2}};q,\bar{q};\mathbb{C}^{2}\right)=1+\sum_{|\mathcal{Y}|}\frac{\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}q^{2}}\right)^{|\mathcal{Y}|}}{\prod_{s\in\mathcal{Y}}\left(1-q^{-l(s)}\bar{q}^{1+a(s)}\right)\left(1-q^{1+l(s)}\bar{q}^{-a(s)}\right)}.$ (25)
@@ 3 Palindromic numerators of bigraded symmetric orbifold Hilbert series and Kostka-Foulkes polynomials
Denoted $K_{\lambda\mu}(q)$, the Kostka-Foulkes polynomials appear in the classical theory of Hall-Littlewood polynomials [8], which are symmetric functions with coefficients depending on a parameter $q$ generalized by the $q,t$ Macdonald polynomials, and are a specialization of their bivariate analogs, the $q,t$ Kostka-Macdonald polynomials $K_{\lambda\mu}(q,t)$. Since the goal of this section involves applications of Schur functions and (transformed) Hall-Littlewood polynomials, we briefly introduce these combinatorial objects before presenting our main result.
@@ 3.1 Hall-Littlewood polynomials
Implicitly introduced in terms of the Hall algebra by Philip Hall and later explicitly defined by D.E. Littlewood, the Hall-Littlewood polynomials provide a generalization for the complete homogeneous symmetric functions as well as the Schur symmetric functions, which will both play a role in our interpretation of the palindromic numerators in terms of the Kostka-Foulkes polynomials. The theory of Hall-Littlewood polynomials and its many connection with algebra, geometry and combinatorics can be found in [8]. In Chapter II for instance, the theory describes the structure of the Hall algebra of finite $\mathfrak{o}$-modules with discrete valuation ring $\mathfrak{o}$, and Chapter VI explains the important role the Hall-Littlewood polynomials play in the representation theory of $GL_{n}\left(\mathbb{F}_{q}\right)$, where $\mathbb{F}_{q}$ is a finite field with $q$ elements.
Beginning with the classical definition of the Hall-Littlewood polynomials [8], for a positive integer $k$, define its $t$-analog and $t$-factorial as
@@
where $\lambda=\left(1^{\alpha_{1}},2^{\alpha_{2}},\ldots\right)$, with $\alpha_{0}$ defined so as to make $\sum_{i}\alpha_{i}=n$, and $x^{\lambda}$ is shorthand for $x_{1}^{\lambda_{1}}x_{2}^{\lambda_{2}}\cdots x_{l}^{\lambda_{l}}$. In the above definition, note that the Hall-Littlewood polynomials generalize the complete homogeneous symmetric functions by letting $t\rightarrow 1$, and the Schur symmetric functions by letting $t\rightarrow 0$.
@@
$Q_{\mu}(X;t)=(1-t)^{\ell(\mu)}P_{\mu}(X;t),\prod_{i}\left[\alpha_{i}\right]_{t}!,$ (28)
@@
$Q_{\lambda}(X;t):=H_{\lambda}\left((1-t)X;t\right),$
@@
$H_{\lambda}(X;t)=Q_{\lambda}\left(\frac{X}{(1-t)};t\right).$
The transformed Macdonald polynomials $H_{\lambda}(X;q,t)$ generlizes the transformed Hall-Littlewood polynomials as, by setting $q=0$, $H_{\mu}(X;0,t)=$ $H_{\mu}(X;t)$.
@@
The conjecture on the positivity of the coefficients in the Kostka-Foulkes polynomials [17], was the prelude of Macdonald’s positivity conjecture, and it’s proof [18] became the inspiration for Macdonald’s positivity theorem.

### v3

@@ 1 Introduction
In [Mvondo-She:2018htn], we introduced and exploited properties of counting formulae for the 1-loop graviton partition function of log gravity [Grumiller:2008qz, Grumiller:2013at] on the thermal $\text{AdS}_{3}$ background originally calculated in [Gaberdiel:2010xv], with the aim of providing a unifying interpretation for both single and multiparticle log sectors. It was shown that the multi-particle states are simply built as a Fock space from products of the single particle states, by reformulating the log partition function in terms of the plethystic exponential (PE), which defines the $n^{\text{th}}$ symmetric product of functions.
Subsequently, we showed in [Mvondo-She:2019vbx] that the moduli space of the log sector is the $S^{n}\left(\mathbb{C}^{2}\right)$ Calabi-Yau orbifold. The modus operandi was to show that the log partition function is a generating function of Hilbert series of the polynomial ring $\mathbb{C}\left[x_{1},\ldots,x_{n},y_{1},\ldots,y_{n}\right]^{S_{n}}$ which is Cohen-Macaulay, and that the numerators of the Hilbert series are palindromic, therefore indicating that the ring being enumerated is Gorenstein, and thus establishing the Calabi-Yau nature of the log sector’s moduli space.
As an expansion in terms of Schur polynomials, i.e Schur symmetric functions expressed as polynomials of power sums, in [Mvondo-She:2021joh] we showed that the log partition function is a $\tau$-function of the KP hierarchy of soliton evolution equations, using the relationship between Schur functions and infinite-dimensional Grassmann manifold studied in [sato1983soliton]. As such, the free energy in (i.e the logarithm of) the log partition function is a solution of the Kadomtsev-Petviashvili (KP) hierarchy that is isomorphic to an infinite-dimensional Grassmann manifold called the Universal (or Sato’s) Grassmann manifold, and the evolution of the log partition function is parametrized by the motion of a point on that Grassmannian.
@@ 2 Background
The aim of this section is to give a succinct expository account of topics that play a central role in our work. Starting with some definitions of symmetric functions, we provide a brief description of the plethystic substitution notation for symmetric functions, which is essential in subsequently introducing Macdonald polynomials, their associated $q,t$-Kostka polynomials, as well as their geometrical interpretation in terms of bigraded $S_{n}$-representations related to the geometry of the Hilbert scheme of $n$ points in the plane.
@@ 2.1 Symmetric functions
A few well-known natural bases for the ring of symmetric functions [macdonald1998symmetric, stanley2011enumerative], all indexed by partitions $\lambda=\left(\lambda_{1},\ldots,\lambda_{k}\right)$ are listed below.
@@
$m_{\lambda}=\sum_{\begin{subarray}{c}i_{a}\neq i_{b}\penalty 10000\ \forall a,b,\\ i_{a}<i_{b}\penalty 10000\ \text{if}\penalty 10000\ \lambda_{a}=\lambda_{b}\end{subarray}}x_{i_{1}}^{\lambda_{1}}\cdots x_{i_{k}}^{\lambda_{k}}.$ (3)
@@
The irreducible representations of the symmetric group $S_{n}$ are also indexed by partitions $\lambda\vdash n$ [macdonald1998symmetric, sagan2013symmetric]. Given an $S_{n}$-module $V$, denote the irreducible $S_{n}$-module by $V_{\lambda}$ such that
@@ 2.3.1 Macdonald polynomials and (q,t)-Kostka Macdonald coefficients
A remarkable new basis for the space of symmetric functions known as the Macdonald polynomials was introduced in [macdonald1988new]. The elements of this basis are denoted $P_{\lambda}\left(X;q,t\right)$ where $\lambda$ is a partition, $X$ is the infinite set of variables $x_{1},x_{2},\ldots$, and $q,t$ are parameters. Their importance stems in part from the fact that by suitable substitutions for the parameters $q$ and $t$, they specialize to many of the well-known bases for the symmetric functions, and one can recover in this manner the Schur functions, the Hall-Littlewood symmetric functions, the Jack symmetric functions, the zonal symmetric functions, the zonal spherical functions, and the elementary and monomial symmetric functions [macdonald1998symmetric].
Given a partition $\lambda$, we define the Young diagram $\mathcal{Y}(\lambda)$ of the partition as the array of lattice squares $\mathcal{Y}(\lambda)=\left\{(i,j)\in\mathbb{N}\times\mathbb{N}:j<\lambda_{i+1}\right\}$, and represent it using the French convention as follows. Given that the parts of $\lambda$ are $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{k}>0$, we let the corresponding Young diagram have $\lambda_{i}$ lattice squares in the $i^{th}$ row (counting from the bottom up), with $n_{\lambda}=\sum_{i=1}^{k}(i-1)\lambda_{i}$, and adopt the Macdonald convention [macdonald1998symmetric] of calling the armarm, legleg, coarmcoarm and colegcoleg of a lattice square $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s$} the parameters $a_{\lambda}$, $l_{\lambda}$, $a^{\prime}_{\lambda}$ and $l^{\prime}_{\lambda}$ giving the number of cells of $\lambda$ that are East, North, West and South of $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s$}, respectively. In this setting, $n_{\lambda}=\sum_{s\in\mathcal{Y}(\lambda)}l(s)$.
$^{\prime} aa$
@@ 2.3.2 Geometry of configuration of $n$ points in the plane
Macdonald conjectured the $K_{\lambda\mu}\left(q,t\right)$ to be polynomials in $q,t$ with positive integer coefficients. In an attempt to prove Macdonald’s conjecture, the so-called modified Macdonald polynomials were introduced in [garsia1993graded] as follows. Let
$H_{\mu}\left(X;q,t\right)=J_{\mu}\left[\frac{X}{1-t};q,t\right]$ (19)
@@
The remarkable connection between the two-parameter symmetric polynomials discovered in [macdonald1988new], and the geometry of configurations of $n$ points in the affine plane $\mathbb{C}^{2}$ can be summarized as follows. In [garsia1993graded], the authors proposed what became known as the $n!$ conjecture, which states that the transformed Macdonald polynomials $\tilde{H}_{\mu}\left(X;q,t\right)$ could be realized as the bigraded characters of certain modules for the diagonal action of $S_{n}$ on two sets of variables. The $n!$ conjecture was proved in [haiman2001hilbert], consequently establishing Macdonald positivity, i.e $K_{\lambda\mu}\left(q,t\right)\in\mathbb{N}[q,t]$, by connecting a conjectural representation-theoretic interpretation of the transformed $q,t$-Kostka coefficients to the algebraic geometry of the Hilbert scheme of $n$ points in the plane, $\text{Hilb}^{n}\left(\mathbb{C}^{2}\right)$, which can be described as the algebraic variety parametrizing ideals $I\subseteq\mathbb{C}\left[x,y\right]$, where $\mathbb{C}\left[x,y\right]$ denotes the coordinate ring of the affine plane, such that dim[$_{\mathbb{C}}\mathbb{C}\left[x,y\right]/I=n$ (i.e the quotient $\mathbb{C}\left[x,y\right]/I$ has dimension $n$ as a vector space over $\mathbb{C}$), or equivalently 0-dimensional subschemes of length $n$ in $\mathbb{C}^{2}$ [haiman2002vanishing, haiman2002combinatorics].
@@ 2.3.3 The log partition
In [Mvondo-She:2019vbx], we showed that the moduli space of the log sector of log gravity is the orbit variety $\left(\mathbb{C}^{2}\right)^{n}/S_{n}$, where $S_{n}$ acts on $\left(\mathbb{C}^{2}\right)^{n}$ by permuting coordinates. A standard procedure when discussing $S^{n}\left(\mathbb{C}^{2}\right)$ is to start by considering ordered $n$-tuples of points in the plane, which we can denote by $P_{1},\ldots,P_{n}\in\mathbb{C}^{2}$. Assigning the points coordinates $x_{1},y_{1},\ldots x_{n},y_{n}$, the space of all $n$-tuples $\left(P_{1},\ldots,P_{n}\right)$ is identified with $\mathbb{C}^{2n}$. The coordinate ring of $\mathbb{C}^{2n}$ is then the polynomial ring $\mathbb{C}\left[\bf{x,y}\right]=\mathbb{C}\left[x_{1},y_{1},\ldots,x_{n},y_{n}\right]$ in $2n$ variables. It follows from the action of the symmetric group $S_{n}$ on $\mathbb{C}^{2n}$ that the subring of invariants $\mathbb{C}\left[\bf{x,y}\right]^{S_{n}}=\mathbb{C}\left[x_{1},y_{1},\ldots,x_{n},y_{n}\right]^{S_{n}}$ is a direct summand of $\mathbb{C}\left[\bf{x,y}\right]$ as an $S_{n}$-module, and also a $\mathbb{C}\left[\bf{x,y}\right]^{S_{n}}$-module. Furthermore, the Hilbert series of $\mathbb{C}\left[\bf{x,y}\right]^{S_{n}}$ can be computed as a doubly graded ring, by degree in the $\bf{x}$ and $\bf{y}$ variables separately.
@@
In addition to the expression (23) formulated in [Mvondo-She:2019vbx], following Macdonald and considering Young diagrams $\mathcal{Y}$ of size $n$, the partition function of the log sector can also be expressed as [^1] Similar expressions have appeared for instance in Eq. (4.5) of [Nakajima:2003pg], [macdonald1998symmetric] or [haiman2002vanishing]..
$Z_{log}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}q^{2}};q,\bar{q};\mathbb{C}^{2}\right)=\sum_{|\mathcal{Y}|}\frac{\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}q^{2}}\right)^{|\mathcal{Y}|}}{\prod_{s\in\mathcal{Y}}\left(1-q^{-l(s)}\bar{q}^{1+a(s)}\right)\left(1-q^{1+l(s)}\bar{q}^{-a(s)}\right)}.$ (25)
@@ 3 Palindromic numerators of bigraded symmetric orbifold Hilbert series and Kostka-Foulkes polynomials
Denoted $K_{\lambda\mu}(q)$, the Kostka-Foulkes polynomials appear in the classical theory of Hall-Littlewood polynomials [macdonald1998symmetric], which are symmetric functions with coefficients depending on a parameter $q$ generalized by the $q,t$ Macdonald polynomials, and are a specialization of their bivariate analogs, the $q,t$ Kostka-Macdonald polynomials $K_{\lambda\mu}(q,t)$. Since the goal of this section involves applications of Schur functions and (transformed) Hall-Littlewood polynomials, we briefly introduce these combinatorial objects before presenting our main result.
@@ 3.1 Hall-Littlewood polynomials
Implicitly introduced in terms of the Hall algebra by Philip Hall and later explicitly defined by D.E. Littlewood, the Hall-Littlewood polynomials provide a generalization for the complete homogeneous symmetric functions as well as the Schur symmetric functions, which will both play a role in our interpretation of the palindromic numerators in terms of the Kostka-Foulkes polynomials. The theory of Hall-Littlewood polynomials and its many connections with algebra, geometry and combinatorics can be found in [macdonald1998symmetric]. In Chapter II for instance, the theory describes the structure of the Hall algebra of finite $\mathfrak{o}$-modules with discrete valuation ring $\mathfrak{o}$, and Chapter VI explains the important role the Hall-Littlewood polynomials play in the representation theory of $GL_{n}\left(\mathbb{F}_{q}\right)$, where $\mathbb{F}_{q}$ is a finite field with $q$ elements.
Beginning with the classical definition of the Hall-Littlewood polynomials [macdonald1998symmetric], for a positive integer $k$, define its $t$-analog and $t$-factorial as
@@
where $\lambda=\left(1^{\alpha_{1}},2^{\alpha_{2}},\ldots\right)$, with $\alpha_{0}$ defined so as to make $\sum_{i}\alpha_{i}=n$, and $x^{\lambda}$ is shorthand for $x_{1}^{\lambda_{1}}x_{2}^{\lambda_{2}}\cdots x_{l}^{\lambda_{l}}$. In the above definition, note that the Hall-Littlewood polynomials generalize the monomial symmetric functions by letting $t\rightarrow 1$, and the Schur symmetric functions by letting $t\rightarrow 0$.
@@
$Q_{\mu}(X;t)=(1-t)^{\ell(\mu)}P_{\mu}(X;t)\prod_{i}\left[\alpha_{i}\right]_{t}!,$ (28)
@@
$Q_{\lambda}[X;t]:=H_{\lambda}\left[(1-t)X;t\right],$
@@
$H_{\lambda}[X;t]=Q_{\lambda}\left[\frac{X}{(1-t)};t\right].$
The transformed Macdonald polynomials $H_{\lambda}(X;q,t)$ generalizes the transformed Hall-Littlewood polynomials as, by setting $q=0$, $H_{\mu}(X;0,t)=$ $H_{\mu}(X;t)$.
@@
The conjecture on the positivity of the coefficients in the Kostka-Foulkes polynomials [foulkes1974survey] was the prelude of Macdonald’s positivity conjecture.